# On the Effectiveness of Methods and Metrics for Explainable AI in Remote Sensing Image Scene Classification

### v1

@@ V Experimental Results
V A Evaluation of the considered explanation metrics
Figure 6: Mean MC scores for the considered explanation metrics: a) scores associated to all RS datasets (DG-ML, FBP-ML and BEN-v2.0); b) scores associated to each individual dataset (DG-ML, FBP-ML, BEN-v2.0). Metrics are categorized into five color-coded groups: Complexity (brown), Faithfulness (blue), Localization (orange), Randomization (green), and Robustness (red).
In Fig. 6(a), we report the mean MC score of each explanation metric, where the scores have been averaged over all three multi‑label RS datasets (BEN-v2.0, DG-ML, and FBP-ML), resulting in a single bar per metric. In contrast, Fig. 6(b) decomposes these results by dataset: for each explanation metric, three separate bars show its mean MC score on DG-ML, FBP-ML and BEN-v2.0 individually. Across both figures, the magnitude of the metrics remains stable, indicating that metric reliability is robust whether computed on the combined dataset or on each dataset in isolation.
From Fig. 6(a) one can see that the metrics from the robustness category are the most reliable for RS image scene classification. In particular, LLE achieves the highest mean MC score of $0.79$, while AS attains the second-highest score of $0.68$. This is because the metrics in this category have fewer limitations compared to other categories and are therefore particularly well-suited for RS image scene classification (see Section III B 2). By observing the figure, one can also see that the localization metrics, RRA and TKI, perform worst with mean MC scores of $0.52$ and $0.51$, respectively. The primary cause of the low performance of localization metrics arises from their methodological constraint: They are unreliable when there is a single LULC class present in $\bm{x}$.
@@ V A Evaluation of the considered explanation metrics
Figure 7: Results of the evaluation of the input test on the DG-ML dataset for the localization metrics: a) Top-K Intersection and b) Relevance Rank Accuracy.
@@
The NR scores indicate consistency over minor perturbations, while the AR scores reflect changes in response to disruptive perturbations. The pattern observed in these scores responds to the methodological flaws of the localization metrics described in Section III C 2. Specifically, when a single LULC class spans the entire image, the explanation coincides exactly with the reference map. Therefore, the metric always yields an output value of $1$. Since the $\text{IAC}_{\text{ER}}$ and $\text{IEC}_{\text{ER}}$ scores measure the metrics consistency over minor perturbations, both scores will consistently assign the highest score of 1. However, as the output is constant, disruptive perturbations cannot change the output of the metric, leading to low $\text{IAC}_{\text{AR}}$ and $\text{IEC}_{\text{AR}}$ values. This flaw is particularly visible in the DG-ML dataset, as it has the lowest average amount of LULC classes per image of the considered datasets.
@@
Furthermore, it is visible in Fig. 6(a) that the randomization metrics also perform comparatively well with the second highest average scores. This is in accordance with the assumption that the metrics of this category perform independently of the data. From the figure, one can see that the MPRT metric achieves a mean MC score of $0.68$, while the RL metric attains a score of $0.60$.
@@
Figure 8: Results of the evaluation of the input test on the DG-ML dataset for the complexity metrics: a) Complexity; and b) Sparseness.
@@
The mean MC values for the metrics CO and SP in the complexity category are $0.63$ and $0.57$, as visualized in Fig. 6(a). Similar to the localization metrics, the complexity metrics also struggle with the characteristic of RS image data in the case that single LULC classes can span large parts of an image, yet only a small region provides the discriminative information necessary for accurate classification. In this case, the metrics tend to attribute higher complexity to the explanation. This deficit is again reflected in the low $\text{IAC}_{\text{AR}}$ and $\text{IEC}_{\text{AR}}$ values, as shown for the CO metric in Fig. 8(a) and the SP metric in Fig. 8(b).
@@
Figure 9: Results of the evaluation of the input test on the BEN-v2.0 dataset for the faithfulness metrics: a) Faithfulness Estimate; and b) Iterative Removal Of Features.
@@
As shown in Fig. 6(a), the mean MC scores for the faithfulness category are $0.56$ and $0.60$ for FE and IROF, respectively. Fig. 9(a) and Fig. 9(b) present the detailed evaluation results for the BEN-v2.0 dataset, which we consider as the perturbation baseline is expected to have a stronger impact for multi-spectral RS images. Similar to the localization metrics, the $\text{IEC}_{\text{AR}}$ score is notably low, with values of 0.28 for FE and 0.01 for IROF. This aligns with methodological findings that faithfulness metrics face similar challenges as localization metrics when an image contains redundant image features for classes that span large parts of the input image.
@@
In such cases, the prediction certainty is not expected to decrease significantly when important features are removed (see Section III A 2). A low $\text{IEC}_{\text{AR}}$ indicates that the metric scores for the perturbed input are consistently lower than for the unperturbed input. However, when the prediction of the model remains unchanged, faithfulness metrics still assign a high score, keeping the score on the unperturbed input high and consequently leading to low $\text{IEC}_{\text{AR}}$ values. Conversely, the $\text{IAC}_{\text{AR}}$ scores for FE and IROF are high, with values of $0.82$ and $1.0$, respectively. This suggests that the distribution of scores remains unchanged even under disruptive perturbations. One possible explanation is a distribution shift induced by OoD samples that occurs due to the perturbation, which prevents the faithfulness scores from effectively capturing the expected changes.
@@
Figure 10: Scores of the IROF metric on the BEN-v2.0 dataset under four perturbation configurations: (i) mean-based LeRF, (ii) mean-based MoRF, (iii) black-based MoRF, and (iv) uniform-based MoRF.
@@
To further investigate how perturbation baselines and removal strategies influence the reliability of faithfulness metrics on multispectral images, we conduct an ablation study using the IROF metric with four configurations: mean-based LeRF, mean-based MoRF, black-based MoRF, and uniform-based MoRF (Fig.10). The results show that combining the mean baseline with LeRF yields the highest reliability (0.75), followed by mean-based MoRF (0.64), black-based MoRF (0.61), and uniform-based MoRF (0.58). As discussed in Section III A 2, removing the most relevant pixels first can fail to capture meaningful changes in prediction certainty, particularly when large image regions contain similar predictive information. In contrast, the LeRF approach preserves critical features until later perturbation steps, leading to a more balanced evaluation. Moreover, employing a mean-value baseline reduces the likelihood of generating OoD samples, further enhancing the reliability of the faithfulness scores.
V B Evaluation of the considered feature attribution methods
@@
TABLE II: Average metric scores obtained by the considered explanation methods for BEN-v2.0
@@
Explanation Method FE IROF AS LLE TKI RRA SP CO RL MPRT
DeepLIFT 0.66 0.59 0.49 1.00 0.49 0.48 0.60 0.30 0.40 0.49
GradCAM 0.73 0.56 0.88 0.99 0.50 0.43 0.46 0.47 0.81 0.67
LIME 0.61 0.54 0.50 0.05 0.49 0.44 0.58 0.46 0.66 0.26
LRP 0.61 0.51 0.71 0.99 0.48 0.45 0.55 0.20 0.23 0.25
Occlusion 0.64 0.48 0.87 0.71 0.49 0.44 0.35 0.52 0.74 0.42
@@
TABLE III: Average metric scores obtained by the considered explanation methods for DG-ML
@@
Explanation Method FE IROF AS LLE TKI RRA SP CO RL MPRT
DeepLIFT 0.67 0.65 0.41 1.00 0.62 0.45 0.71 0.48 0.46 0.66
GradCAM 0.65 0.65 0.68 0.96 0.74 0.48 0.43 0.42 0.83 0.49
LIME 0.53 0.66 0.66 0.56 0.64 0.44 0.52 0.50 0.72 0.23
LRP 0.69 0.66 0.61 0.99 0.64 0.47 0.54 0.26 0.22 0.42
Occlusion 0.75 0.68 0.69 0.43 0.69 0.47 0.31 0.40 0.82 0.34
@@
TABLE IV: Average metric scores obtained by the considered explanation methods for FBP-ML
@@
Explanation Method FE IROF AS LLE TKI RRA SP CO RL MPRT
DeepLIFT 0.75 0.59 0.45 1.00 0.46 0.49 0.70 0.43 0.29 0.69
GradCAM 0.75 0.62 0.75 0.99 0.48 0.50 0.57 0.85 0.71 0.46
LIME 0.65 0.55 0.56 0.41 0.45 0.45 0.55 0.51 0.67 0.16
LRP 0.69 0.58 0.53 0.99 0.43 0.47 0.65 0.28 0.29 0.26
Occlusion 0.81 0.58 0.69 0.81 0.46 0.48 0.50 0.48 0.73 0.25
@@
Localization is evaluated using the TKI and RRA metrics. Across all datasets and metrics, except for RRA on BEN-v2.0, GradCAM performs best. The variance across the scores for the different datasets and methods remains low, except for the TKI metric in the DG-ML dataset, which exhibits higher variability. Complexity is assessed using the SP and CO metrics. For the SP metric, DeepLIFT achieves the highest performance across all datasets. The results for the CO metric vary across datasets; averaged over all datasets, GradCAM yields the best performance with an average score of 0.58. The random baseline performs poorly, as expected, given that the CO metric is defined as the entropy of the explanation, which reaches its maximum when the explanation is uniformly distributed. Randomization is quantified using the RL and MPRT metrics. The random baseline achieves the highest RL score across all datasets, with an average performance of 0.97, aligning with the findings of [66], which demonstrate that randomized attribution maps can achieve high scores in top-down randomization metrics. While this outcome exposes a limitation of the RL metric, it remains valuable in identifying explanation methods that are largely independent of the model and therefore unreliable. For the MPRT metric, GradCAM achieves the best result on the BEN-v2.0 dataset, while DeepLIFT performs best on DG-ML and FBP-ML.
@@
Fig. 12(a) displays the overall average performance of the evaluated explanation methods across all considered metrics. The results show that GradCAM achieves the highest score ($0.65$), followed by DeepLIFT ($0.58$) and Occlusion ($0.57$). LRP and LIME achieve slightly lower scores of $0.52$ and $0.50$, respectively, while the Random Baseline performs worst at $0.46$. In contrast, Fig. 12(b) presents the weighted average scores of the most reliable metrics. Here, GradCAM again performs best ($0.66$), but DeepLIFT achieves a higher relative score ($0.63$), surpassing Occlusion ($0.52$). Notably, LRP improves in ranking, reaching a score of $0.53$, while LIME and the Random Baseline exhibit a more pronounced decline, with scores of $0.43$ and $0.37$, respectively. Comparing both figures, GradCAM consistently outperforms other methods, maintaining the highest rank across both evaluation approaches. However, the usage of more reliable metrics in Fig. 12(b) results in an increased performance gap between methods, with DeepLIFT improving its relative standing and LIME and the Random Baseline exhibiting a sharper decline. This shows that using more reliable metrics leads to a more accurate evaluation and confirms the robustness of GradCAM. The larger reduction in Random Baseline performance indicates that the most reliable metrics for RS image scene classification effectively penalize random explanations.
Figure 11: Comparison of metric averages across datasets. Each metric is color-coded by category.
Figure 12: Comparison of explanation methods based on evaluation metrics. (a): normalized average of all metrics across all RS datasets; (b): normalized average of the best-performing metric (CO, IROF, LLE, RRA, MPRT) for each category across all RS datasets.
@@
VI Conclusion and Guidelines
@@
Our methodological analysis shows that perturbation‐based approaches, such as Occlusion and LIME, are critically dependent on the choice of perturbation baseline. In RS scenes, replacing spectral content with a fixed value can generate OoD samples. Furthermore, the redundancy of features across large regions hinders the perturbation function to significantly decrease the prediction of a model, which lowers the reliability of perturbation-based methods. Gradient‐propagation methods like GradCAM are limited in their ability to distinguish between multiple instances of the same object in a scene. Relevancy-propagating methods like LRP tend to underestimate the contribution of LULC classes that span larger areas of a scene. Regarding evaluation metrics, our findings indicate that faithfulness metrics inherit the limitations of perturbation‐based methods and need to be parametrized accordingly. In contrast, robustness metrics demonstrate greater stability. Localization metrics require accurate pixel-level reference maps. Furthermore, localization and complexity metrics are unreliable when a single class spans a large part of the scene. Randomization metrics are often independent of the data. To support our methodological findings, we conducted two sets of experiments on three RS datasets: BEN-v2.0, DG-ML, and FBP-ML. In the first set of experiments, we meta-evaluated ten explanation metrics to assess their reliability. We observe that metrics from the robustness and randomization categories performed well. Conversely, as expected from the methodological analysis, metrics from the faithfulness and localization categories are more unreliable. In the second set of experiments, we use the most reliable explanation metrics to determine which explanation methods are best suited for RS image scene classification. Among the evaluated methods, GradCAM emerged as the most robust feature attribution method, demonstrating relatively stable performance across different datasets and evaluation metrics. However, it is important to note that no single method performed optimally across all evaluation criteria.
@@
    We suggest using localization metrics only when accurate pixel-level reference maps are available. In particular, we recommend using more advanced localization metrics that also account for the size of the target class. Complexity metrics should be interpreted carefully, as they depend strongly on the spatial extent of the target class. Similarly to localization metrics, complexity metrics could also be modified to account for the target class size.
@@
    We recommend using GradCAM as a general baseline due to its versatility and ease of interpretation. Among the evaluated methods, it consistently achieves high scores across multiple metric categories.
As a final remark, we would like to note that, due to the lack of ground truth on explanations, it remains not possible to fully validate an explanation metric [24]. The quantifiable reliability of a metric does not necessarily indicate its validity [66]. Thus, each evaluation must be critically examined. Future work should focus on refining explanation metrics specifically designed for RS image scene classification. In particular, developing perturbation strategies that better accommodate spectral variability and feature redundancy may enhance the reliability of faithfulness metrics and, by extension, perturbation-based explanation methods. Although our study concentrated on feature attribution methods commonly used in RS (while being originally designed for natural images), a promising direction for future research is to develop and evaluate explanation techniques tailored specifically to the characteristics of RS images.

### v3

@@ V Experimental Results
V.A Evaluation of the considered explanation metrics
In Fig. 6(a), we report the mean MC score of each explanation metric, where the scores have been averaged over all three multi‑label RS datasets (BEN-v2.0, DG-ML, and FBP-ML), resulting in a single bar per metric. In contrast, Fig. 6(b) decomposes these results by dataset. For each explanation metric, three separate bars show its mean MC score on DG-ML, FBP-ML, and BEN-v2.0 individually. Across both figures, the magnitude of the metrics remains stable, indicating that metric reliability is robust whether computed on the combined dataset or on each dataset in isolation.
From Fig. 6(a), one can see that the metrics from the robustness category are the most reliable for RS image scene classification. In particular, LLE achieves the highest mean MC score of $0.79$, while AS attains the second-highest score of $0.68$. This is because the metrics in this category have fewer limitations compared to other categories and are therefore particularly well-suited for RS image scene classification (see Section III.B.2). By observing the figure, one can also see that the localization metrics, RRA and TKI, perform worst with mean MC scores of $0.52$ and $0.51$, respectively. The primary cause of the low performance of localization metrics arises from their methodological constraint, which is that they are unreliable when there is a single LULC class present in $\bm{x}$.
@@ V.A Evaluation of the considered explanation metrics
Figure 7: Results of the evaluation protocol on the DG-ML dataset [46] for the localization metrics: a) Top-K Intersection and b) Relevance Rank Accuracy.
@@
The NR scores indicate consistency over minor perturbations, while the AR scores reflect changes in response to disruptive perturbations. The pattern observed in these scores responds to the methodological flaws of the localization metrics described in Section III.C.2. Specifically, when a single LULC class spans the entire image, the explanation coincides exactly with the reference map. Therefore, the metric always yields an output value of $1$. Since the $\text{IAC}_{\text{ER}}$ and $\text{IEC}_{\text{ER}}$ scores measure the metrics consistency over minor perturbations, both scores consistently assign the highest score of 1. However, as the output is constant, disruptive perturbations cannot change the output of the metric, leading to low $\text{IAC}_{\text{AR}}$ and $\text{IEC}_{\text{AR}}$ values. This flaw is particularly visible in the DG-ML dataset, as it has the lowest average amount of LULC classes per image of the considered datasets.
@@
Furthermore, it is visible in Fig. 6(a) that the randomization metrics also perform comparatively well with the second-highest average scores. This is in accordance with the assumption that the metrics of this category perform independently of the data. From the figure, one can see that the MPRT metric achieves a mean MC score of $0.68$, while the RL metric attains a score of $0.60$.
@@
Figure 8: Results of the evaluation protocol on the DG-ML dataset [46] for the complexity metrics: a) Complexity; and b) Sparseness.
@@
The mean MC values for the metrics CO and SP in the complexity category are $0.63$ and $0.57$, as visualized in Fig. 6(a). Similar to the localization metrics, the complexity metrics also struggle with the characteristic of RS image data. Single LULC classes can span large parts of an image, yet only a small region provides the discriminative information necessary for accurate classification. In this case, the metrics tend to attribute higher complexity to the explanation. This deficit is again reflected in the low $\text{IAC}_{\text{AR}}$ and $\text{IEC}_{\text{AR}}$ values, as shown for the CO metric in Fig. 8(a) and the SP metric in Fig. 8(b).
@@
Figure 9: Results of the evaluation of the evaluation protocol on the BEN-v2.0 dataset [60] for the faithfulness metrics: a) Faithfulness Estimate; and b) Iterative Removal Of Features.
@@
As shown in Fig. 6(a), the mean MC scores for the faithfulness category are $0.56$ and $0.60$ for FE and IROF, respectively. Fig. 9(a) and Fig. 9(b) present the detailed evaluation results for the BEN-v2.0 dataset, which we consider as the perturbation baseline is expected to have a stronger impact for multispectral RS images. Similar to the localization metrics, the $\text{IEC}_{\text{AR}}$ score is notably low with values of 0.28 for FE and 0.01 for IROF. This aligns with methodological findings that faithfulness metrics face similar challenges as localization metrics when an image contains redundant image features for classes that span large parts of the input image.
@@
In such cases, the prediction certainty is not expected to decrease significantly when important features are removed (see Section III.A.2). A low $\text{IEC}_{\text{AR}}$ indicates that the metric scores for the perturbed input are consistently lower than for the unperturbed input. However, when the prediction of the model remains unchanged, faithfulness metrics still assign a high score, keeping the score on the unperturbed input high and consequently leading to low $\text{IEC}_{\text{AR}}$ values. On the contrary, the $\text{IAC}_{\text{AR}}$ scores for FE and IROF are high, with values of $0.82$ and $1.0$, respectively. This suggests that the distribution of scores remains unchanged even under disruptive perturbations. One possible explanation is a distribution shift caused by OoD samples introduced through the perturbation, which prevents the faithfulness scores from effectively capturing the expected changes.
@@
Figure 10: MC Scores of the IROF metric on the BEN-v2.0 dataset [60] under four perturbation configurations: (i) mean-based LeRF, (ii) mean-based MoRF, (iii) black-based MoRF, and (iv) uniform-based MoRF.
@@
To further investigate how perturbation baselines and removal strategies influence the reliability of faithfulness metrics on multispectral images, we conduct an ablation study using the IROF metric with four configurations: i) mean-based LeRF; ii) mean-based MoRF; iii) black-based MoRF; and iv) uniform-based MoRF (Fig.10). The results show that combining the mean baseline with LeRF yields the highest reliability (0.75), followed by mean-based MoRF (0.64), black-based MoRF (0.61), and uniform-based MoRF (0.58). As discussed in Section III.A.2, removing the most relevant pixels first can fail to capture meaningful changes in prediction certainty, particularly when large image regions contain similar predictive information. In contrast, the LeRF approach preserves critical features until later perturbation steps, leading to a more balanced evaluation. Moreover, employing a mean-value baseline reduces the likelihood of generating OoD samples, further enhancing the reliability of the faithfulness scores.
V.B Evaluation of the considered feature attribution methods
@@
TABLE II: Average metric scores obtained by the considered explanation methods for BEN-v2.0 [60].
@@
Explanation Method FE [28] IROF [29] AS [30] LLE [28] TKI [32] RRA [31] SP [33] CO [34] RL [36] MPRT [59]
DeepLIFT [12] 0.66 0.59 0.49 1.00 0.49 0.48 0.60 0.30 0.40 0.49
GradCAM [13] 0.73 0.56 0.88 0.99 0.50 0.43 0.46 0.47 0.81 0.67
LIME [10] 0.61 0.54 0.50 0.05 0.49 0.44 0.58 0.46 0.66 0.26
LRP [11] 0.61 0.51 0.71 0.99 0.48 0.45 0.55 0.20 0.23 0.25
Occlusion [9] 0.64 0.48 0.87 0.71 0.49 0.44 0.35 0.52 0.74 0.42
@@
TABLE III: Average metric scores obtained by the considered explanation methods for DG-ML [46].
@@
Explanation Method FE [28] IROF [29] AS [30] LLE [28] TKI [32] RRA [31] SP [33] CO [34] RL [36] MPRT [59]
DeepLIFT [12] 0.67 0.65 0.41 1.00 0.62 0.45 0.71 0.48 0.46 0.66
GradCAM [13] 0.65 0.65 0.68 0.96 0.74 0.48 0.43 0.42 0.83 0.49
LIME [10] 0.53 0.66 0.66 0.56 0.64 0.44 0.52 0.50 0.72 0.23
LRP [11] 0.69 0.66 0.61 0.99 0.64 0.47 0.54 0.26 0.22 0.42
Occlusion [9] 0.75 0.68 0.69 0.43 0.69 0.47 0.31 0.40 0.82 0.34
@@
TABLE IV: Average metric scores obtained by the considered explanation methods for FBP-ML [62].
@@
Explanation Method FE [28] IROF [29] AS [30] LLE [28] TKI [32] RRA [31] SP [33] CO [34] RL [36] MPRT [59]
DeepLIFT [12] 0.75 0.59 0.45 1.00 0.46 0.49 0.70 0.43 0.29 0.69
GradCAM [13] 0.75 0.62 0.75 0.99 0.48 0.50 0.57 0.85 0.71 0.46
LIME [10] 0.65 0.55 0.56 0.41 0.45 0.45 0.55 0.51 0.67 0.16
LRP [11] 0.69 0.58 0.53 0.99 0.43 0.47 0.65 0.28 0.29 0.26
Occlusion [9] 0.81 0.58 0.69 0.81 0.46 0.48 0.50 0.48 0.73 0.25
@@
Localization is evaluated using the TKI and RRA metrics. Across all datasets and metrics, except for RRA on BEN-v2.0, GradCAM performs best. The variance across the scores for the different datasets and methods remains low, except for the TKI metric in the DG-ML dataset, which exhibits higher variability. Complexity is assessed using the SP and CO metrics. For the SP metric, DeepLIFT achieves the highest performance across all datasets. The results for the CO metric vary across datasets; averaged over all datasets, GradCAM yields the best performance with an average score of 0.58. The random baseline performs poorly, as expected, given that the CO metric is defined as the entropy of the explanation, which reaches its maximum when the explanation is uniformly distributed. Randomization is quantified using the RL and MPRT metrics. The random baseline achieves the highest RL score across all datasets, with an average performance of 0.97. This aligns with the findings of [67], which demonstrate that randomized attribution maps can achieve high scores in top-down randomization metrics. While this outcome exposes a limitation of the RL metric, it remains valuable in identifying explanation methods that are largely independent of the model and therefore unreliable. For the MPRT metric, GradCAM achieves the best result on the BEN-v2.0 dataset, while DeepLIFT performs best on DG-ML and FBP-ML.
@@
Fig. 12(a) and Fig. 12(b) show a comparison of the overall and weighted average performances of the evaluated explanation methods across all considered metrics. In both evaluations, GradCAM consistently achieves the highest scores. When weighting by the most reliable metrics, the relative differences between methods become more pronounced. In detail, DeepLIFT improves its standing, while LIME and the Random Baseline show a substantial decline. This indicates that reliable metrics provide a sharper separation between strong and weak methods, effectively penalizing random explanations and demonstrating the reliability of GradCAM.
Figure 11: Average performance of explanation metrics for each considered dataset, grouped by evaluation category. Each bar corresponds to one metric, with colors indicating the metric category: brown for complexity, blue for faithfulness, orange for localization, green for randomization, and red for robustness.
Figure 12: Comparison of explanation methods based on evaluation metrics. a): normalized average of all the considered metrics across all the considered RS datasets; b): normalized average of the best-performing metric (CO, IROF, LLE, RRA, MPRT) for each category across all considered RS datasets.
@@
V.C Evaluation of the computational efficiency
TABLE V: Average computation time (in seconds) for metrics and explanation methods on the DG-ML dataset [46].
Group Name Time (seconds)
Explanation Methods DeepLIFT [12] 0.26
GradCAM [13] 0.12
LIME [10] 28.63
LRP [11] 0.38
Occlusion [9] 2.04
Random baseline 0.04
Explanation Metrics FE [28] 3.39
IROF [29] 13.26
AS [30] 12.15
LLE [28] 11.97
TKI [32] 0.10
RRA [31] 0.16
SP [33] 0.06
CO [34] 0.06
RL [36] 0.03
MPRT [59] 11.10
In this subsection, we analyze the computational efficiency of the considered metrics and explanation methods for RS scene classification. Table V reports the average runtime required in seconds to compute each explanation method or metric on the DG-ML dataset.
From the table, one can observe that the average computation time required for the relevancy-propagation-based and gradient-propagation-based methods, such as GradCAM, DeepLIFT, and LRP, is below one second. This is because these methods explain the prediction in a single forward pass. In contrast, perturbation-based approaches such as LIME and Occlusion require substantially more computation, as they require multiple forward passes. The runtime of LIME is particularly high (28.63 seconds) due to the need to evaluate perturbed samples and fit a surrogate model. However, it is worth noting that the efficiency of perturbation-based methods depends mostly on their parametrization, as increasing the number of perturbations improves the reliability of the results but proportionally increases computation time.
For explanation metrics, a similar behavior can be observed from the table. Perturbation-based metrics such as FE, IROF, AS, LLE, and MPRT require several seconds due to repeated recomputation of explanations under perturbations. In contrast, metrics from localization, complexity, and randomization categories, including TKI, RRA, SP, CO, and RL, are computed in closed form from a single attribution and thus achieve runtimes below 0.2 seconds. These results highlight that the computational efficiency of explanation metrics is largely dictated by whether perturbation is required.
VI Conclusion and Discussion
@@
Our methodological analysis shows that perturbation‐based approaches, such as Occlusion and LIME, are critically dependent on the choice of perturbation baseline. In RS scenes, replacing spectral content with a fixed value can generate OoD samples. Furthermore, the redundancy of features across large regions hinders the perturbation function to significantly decrease the prediction of a model, which lowers the reliability of perturbation-based methods. Gradient‐propagation methods like GradCAM are limited in their ability to distinguish between multiple instances of the same object in a scene. Relevancy-propagating methods like LRP tend to underestimate the contribution of LULC classes that span larger areas of a scene. Regarding evaluation metrics, our findings indicate that faithfulness metrics inherit the limitations of perturbation‐based methods and need to be parametrized accordingly. In contrast, robustness metrics demonstrate greater stability. Localization metrics require accurate pixel-level reference maps. Furthermore, localization and complexity metrics are unreliable when a single class spans a large part of the scene. Randomization metrics are often independent of the data. To support our methodological findings, we conducted two sets of experiments on three RS datasets: BEN-v2.0 [60], DG-ML [46], and FBP-ML [62]. In the first set of experiments, we meta-evaluated ten explanation metrics to assess their reliability. We observe that metrics from the robustness and randomization categories are reliable. As expected from the methodological analysis, metrics from the faithfulness and localization categories are less reliable. In the second set of experiments, we use the most reliable explanation metrics to determine which explanation methods are best suited for RS image scene classification. Among the evaluated methods, GradCAM emerges as the most robust feature attribution method, demonstrating relatively stable performance across different datasets and evaluation metrics. However, it is important to note that no single method has performed optimally across all evaluation criteria.
@@
    We suggest using localization metrics only when accurate pixel-level reference maps are available. In particular, we recommend using more advanced localization metrics that also account for the size of the target class. Complexity metrics should be interpreted carefully, as they depend strongly on the spatial extent of the target class. Similar to localization metrics, complexity metrics could also be modified to account for the target class size.
@@
    We recommend using GradCAM as a general baseline due to its versatility, speed, and ease of interpretation. Among the evaluated methods, it consistently achieves high scores across multiple metric categories.
Beyond methodological contributions, our findings have direct implications for operational applications. For instance, reliable explanation methods are crucial in maritime monitoring, where RS scene classification is used for maritime surveillance and decision support. Moreover, reliable attribution maps can serve as pseudo labels in tasks such as tree species classification [68].
As a final remark, we emphasize that the evaluation protocol can not measure the intrinsic validity of explanation metrics, since no ground truth for explanations exists [24]. It can probe the reliability and comparative behavior of metrics, yet reliability alone does not imply validity [67]. Evaluation outcomes should therefore be regarded as indicative rather than definitive, and the theoretical foundations of the metrics also have to be taken into account.
As a future work, we plan to refine explanation metrics specifically designed for RS image scene classification. In particular, we want to develop perturbation strategies that better accommodate spectral variability and feature redundancy, which may enhance the reliability of faithfulness metrics and, by extension, perturbation-based explanation methods. Although our study concentrated on feature attribution methods commonly used in RS (while being originally designed for natural images), a promising direction for future research is to develop explanation methods and metrics tailored specifically to the characteristics of RS images. Another important direction is to evaluate more recent explanation approaches, such as concept-based XAI, and to extend the evaluation framework accordingly.